OpenStudy (anonymous):

Can i get some help with this, Please? Given cos theta=24/25 and is located in Q3, Evaluate the rest of the trigonometric functions: 1. sin theta= 2. tan theta= 3. cot theta= 4. sec theta= 5. csc theta=

jimthompson5910 (jim_thompson5910):

sin^2 + cos^2 = 1 sin^2 + (24/25)^2 = 1 sin^2 + 576/625 = 1 sin^2 = 1 - 576/625 sin^2 = 49/625 sin = sqrt(49/625) sin = 7/25 Since we're in Q3, sine theta is negative, so sin theta = -7/25 So when cos theta=24/25, sin theta = -7/25

OpenStudy (anonymous):

eh, that looks confusing. So for tan theta, would i start off as tan^2+cos^2

jimthompson5910 (jim_thompson5910):

tan = sin/cos

jimthompson5910 (jim_thompson5910):

that's one of the definitions of tangent

OpenStudy (anonymous):

oh yeah, i remember that. So what do i do next?

jimthompson5910 (jim_thompson5910):

tan = sin/cos tan = (-7/25)/(-24/25) tan = (-7/25)*(-25/24) tan = 7/24 Note: cosine should be negative as well in Q3...not sure why they made it positive

OpenStudy (anonymous):

So is there a reason why you swapped (-24/25) to (-25/24) or does it matter?

jimthompson5910 (jim_thompson5910):

(a/b) divided by (c/d) is the same as (a/b) times (d/c)

jimthompson5910 (jim_thompson5910):

when you divide fractions, flip the second fraction and multiply

OpenStudy (anonymous):

Oh ok. So im a little familiar with tan, cos, and sin. But not with cot,csc, and sec.

jimthompson5910 (jim_thompson5910):

cot = 1/tan csc = 1/sin sec = 1/cos all by definition

jimthompson5910 (jim_thompson5910):

so if tan = a/b, then cot = b/a (you flip the fraction...ie find the reciprocal) In this case, tan = 7/24 so cot = 24/7

OpenStudy (anonymous):

so times cot by cos?

jimthompson5910 (jim_thompson5910):

no, you take whatever tan is and you flip it to get cot

OpenStudy (anonymous):

oh, pretty easy.

jimthompson5910 (jim_thompson5910):

yeah it's not too bad once you get the hang of it

jimthompson5910 (jim_thompson5910):

the trick is remembering what cot, csc and sec stand for (and what they equal)

OpenStudy (anonymous):

so sec=1/cos so sec=1/(24/25)?

jimthompson5910 (jim_thompson5910):

yes, but cosine should be negative

jimthompson5910 (jim_thompson5910):

so sec = 1/(-24/25) ----> sec = -25/24

OpenStudy (anonymous):

and the last one is csc theta= what does csc even mean? Like tan=tangent.

jimthompson5910 (jim_thompson5910):

cot = cotangent csc = cosecant sec = secant

OpenStudy (anonymous):

So for csc... 1/(-7/25) ?

jimthompson5910 (jim_thompson5910):

yes, which becomes csc = -25/7

OpenStudy (anonymous):

Your so much help. Thank you so so much.

jimthompson5910 (jim_thompson5910):

you're very welcome